# UnlearnCanvas: A Stylized Image Dataset to Benchmark Machine Unlearning for Diffusion Models

### v2

@@ B Additional Machine Unlearning Experiment Results
Comparison to the most relevant dataset WikiArt.
We first make a comprehensive comparison between UnlearnCanvas and WikiArt on their key properties in Tab. A1. As we can see, UnlearnCanvas out performs WikiArt from the several perspectives. First, UnlearnCanvas contain images of much higher resolution (15M vs. 2M). This in particular benefits the training associated with DMs, where larger images in general help the model captures more details, such as the stylistic strokes. Second, the images in UnlearnCanvas are style-wise supervised. This means, for each seed image, a stylized counterpart can be find in each style class. This is beneficial for other tasks than MU, such as style transfer, which can provide a ground truth image for precise and robust evaluation. This will be detailed in Appx. C. Third, UnlearnCanvas enjoys high intra-style stylistic consistency and inter-style stylistic distinctiveness. This means the images of the same label share much more stylistic similarity, which is easier for a deep learning model to learn and to classify. This benefits can also be reflected by the training performance Appx. A.1 using UnlearnCanvas and WikiArt. The results are reported in Tab. 2 and Tab. A2, respectively. As we can see, the classifier is much more easily trained on UnlearnCanvas, justifying the higher discernible features within each style in UnlearnCanvas.
Table A2: Style classification results of a ViT-Large (Dosovitskiy et al., 2020) as a style classifier trained on UnlearnCanvas. After convergence, the classifier is tested on the test set and the image set generated by SD v1.5 finetuned on UnlearnCanvas.
@@
A fine-grained comparison per unlearning target: ESD vs. Salun.
Following the analysis of ESD, we next turn our focus to a comparative analysis with Salun, a method that demonstrated a better balance between unlearning and retaining according to Table 3. A similar accuracy heatmap for Salun is presented in Figure A1. Compared to ESD, Salun exhibits more consistent performance across various unlearning scenarios, as indicated by the more uniform color distribution in the heatmap. This also suggests enhanced retainability. However, it is noticeable that Salun does not reach the same level of UA (Unlearning Accuracy) as ESD, as evidenced by the darker diagonal values in Figure A1. This observation reinforces the existence of a trade-off between unlearning effectiveness and retainability in the visual generative MU task, a phenomenon paralleled in other tasks such as classification.
@@
C Broader Use Cases of UnlearnCanvas
@@
C.1 Benchmarking Style Transfer using UnlearnCanvas
@@
Style transfer is a long-standing topic and focuses on transferring the artistic style from one style image (also known as the reference image) $\mathbf{x}_{s}$ to a target content image $\mathbf{x}_{c}$. The most recent methods typically employ a neural network, denoted by ${\bm{\theta}}_{s}$, to extract style features and perform stylization in a single inference step, expressed as $\hat{\mathbf{x}}_{\mathrm{o}}=f_{{\bm{\theta}}_{s}}(\mathbf{x}_{s},\mathbf{x}_{c})$. Current state-of-the-art (SOTA) style transfer techniques exhibit remarkable generalization capabilities, successfully transferring styles not encountered during training and not requiring any further back-propagations. Figure 1 illustrates the pipeline of this task. Most existing literature (Liu et al., 2021; Huang & Belongie, 2017; Li et al., 2019; Chen & Schmidt, 2016; Park & Lee, 2019; An et al., 2021) utilize WikiArt (Phillips & Mackintosh, 2011) to provide different styles for training the stylization network ${\bm{\theta}}_{s}$. During the evaluation, the validation set of WikiArt will serve as the style (reference) images, together with the content images from the COCO dataset (Lin et al., 2014), to form a test bed for style transfer and style learning methods. However, such a evaluation scheme has some inherent limitations, which will be detailed below.
@@
Although the task of style transfer has been widely studied, its evaluations are still based on very limited quantitative or even sorely qualitative assessments, potentially leading to incomplete and inaccurate assessments (Cai et al., 2023). Upon examining the evaluation pipelines of over $10$ state-of-the-art (SOTA) style transfer methods, three significant challenges are identified within the current widely accepted evaluation frameworks. $\bullet$ Challenge I (C1): The lack of the ground truth images for style similarity evaluation. Unlike other vision tasks like classification (Singh & Singh, 2020), detection (Zou et al., 2023), and segmentation (Minaee et al., 2021), one of the key shortcomings of the current style transfer evaluation lies in the lack of the ground truth images $\mathbf{x}_{g}$ for the given reference style image $\mathbf{x}_{s}$ and the content image $\mathbf{x}_{c}$. Consequently, existing evaluation metrics, such as the style loss $\ell_{\mathrm{style}}$ (Gatys et al., 2015a, b), has to be calculated with the reference style image $\mathbf{x}_{s}$ as the ground truth $\mathbf{x}_{g}$, namely $\ell_{\mathrm{style}}(\mathbf{x}_{s},\hat{\mathbf{x}}_{o})$, rather than directly using the ground truth $\ell_{\mathrm{style}}(\mathbf{x}_{o},{\mathbf{x}}_{g})$. Obviously, such a indirect evaluation may lead to inaccurate results due to the different contents held in $\mathbf{x}_{s}$ and $\mathbf{x}_{o}$. Existing work has demonstrated that such evaluation metrics can lead to very different conclusion from that of the user study (An et al., 2021). Therefore, the creation of a supervised dataset with ground truth stylized images for every content image under each style is a timely remedy. $\bullet$ Challenge II (C2): The lack of algorithm stability evaluation against varied reference images $\mathbf{x}_{s}$. The evaluation of algorithm stability in style transfer has often been neglected. This assessment requires consistent performance of a method on a content image $\mathbf{x}_{c}$ across different style reference images ${\mathbf{x}_{s}}$ representing the same target artistic style. Ideally, an algorithm should maintain uniform quality across various references within a style, avoiding significant performance variations. Current challenges in such evaluations stem from the lack of reference sets that exhibit high stylistic consistency within each style. Figure 4 showcases examples from the widely-used WikiArt dataset. Despite sharing the same artistic label, images within a row show considerable divergence in visual appearance and style. Consequently, using these images as style references can lead to stable algorithms producing stylistically varied outputs, leading to misleading assessment results. Therefore, creating a dataset with high stylistic uniformity within each style category and clear differentiation between styles is crucial for accurate measurement of algorithm stability.
@@
To revolutionize the evaluation framework with UnlearnCanvas, it is crucial to first understand the components that form a comprehensive evaluation pipeline, ensuring a thorough and impartial assessment of performance. In the realm of style transfer, three pivotal metrics stand paramount, as depicted in Figure,A3:
@@
① Style Loss: Utilizes a feature map-based style loss (Deng et al., 2022) to quantify the stylistic dissimilarity between image pairs, effectively representing the inverse of style authenticity. ② Content Loss: Employs a VGG-based, feature-map content loss (Deng et al., 2022; Zhang et al., 2022) to measure the visual dissimilarity between the reference and generated images, essentially mirroring content integrity. ③ Averaged Standard Deviation (STD): Computes the average STD of style and content loss w.r.t. the same reference image, reflecting algorithm stability.
@@
Following this evaluation pipeline, we scrutinized $9$ prominent style transfer methods, including SaNet (Park & Lee, 2019), Mcc (Deng et al., 2021), Mast (Huo et al., 2021), ArtFlow with its two variants (AF-AdaIn and AF-WCT) (An et al., 2021), IE-Contrast (Chen et al., 2021), Cast (Zhang et al., 2022), StyTr2 (Deng et al., 2022), and BLIP (Li et al., 2023).
@@
C.2 Broader Use Cases of UnlearnCanvas

### v4

@@ C A Detailed Comparison between UnlearnCanvas and WikiArt
Second, UnlearnCanvas surpasses WikiArt in terms of both intra-style coherence and inter-style distinctiveness, as illustrated in Fig. A1, where images labeled with ‘Van Gogh Style’ from both datasets are compared. In UnlearnCanvas, the images exhibit high stylistic consistency, while WikiArt lacks the necessary clarity for precise assessment. This will hamper the MU evaluation as discussed in the challenges (C2) and (C3). This benefits can also be reflected by the training performance using UnlearnCanvas and WikiArt. The results are reported in Tab. A3 and Tab. A4, respectively. As we can see, the classifier is much more easily trained on UnlearnCanvas, justifying the higher discernible features within each style in UnlearnCanvas.
Table A3: Art style reproduction quality using SD v1.5 and SD v2.0 finetuned on WikiArt. Images are generated with the prompt “A painting in artist style”, where artist refers to those included in WikiArt. The test accuracy on DM-generated images and original WikiArt test images is reported using the style classifier finetuned from the pretrained ViT-L/16 on WikiArt.
Image
Source
Images by
SD v1.5
Images by
SD v2.0
WikiArt
Test Set
Accuracy 41.2% 56.7% 85.4%
Table A4: Style classification results of a ViT-Large [68] as a style classifier trained on UnlearnCanvas. After convergence, the classifier is tested on the test set and the image set generated by SD v1.5 finetuned on UnlearnCanvas.
@@
Third, the images in UnlearnCanvas are style-wise supervised. For each seed image, a stylized counterpart can be find in each style class. This is beneficial for tasks other than unlearning for text-to-image task, such as image editing, image stylization, and style transfer, which can provide a ground truth image for precise and robust evaluation. This will be detailed in Appx. F.
D Additional Experiment Results for Unlearning Evaluation
D.1 Visualization of the Style and Object Unlearning Performance
To make a more direct comparison among different MU methods reported in Tab. 1, the results are visualized in the radar chart Fig. A2. This figure illustrates that no method dominates across all assessment dimensions. This underscores the complexity of unlearning in generative models and the need for further improvement.
Figure A2: Performance visualization for various unlearning methods as summarized in Table 2. For UA, IRA, and CRA, the results are averaged over the style and object unlearning scenarios. Other metrics undertake the inverse operation as a smaller values represent better performance. Results are normalized to $0\%\sim 100\%$ per metric.
D.2 A Fine-Grained Comparison per Unlearning Target: ESD vs. Salun
Following the analysis of ESD in Fig. 6, we next turn our focus to a comparative analysis with Salun, a method that demonstrated a better balance between unlearning and retaining according to Tab. 2. A similar accuracy heatmap for Salun is presented in Fig. A3. Compared to ESD, Salun exhibits more consistent performance across various unlearning scenarios, as indicated by the more uniform color distribution in the heatmap. This also suggests enhanced retainability. However, it is noticeable that Salun does not reach the same level of UA (Unlearning Accuracy) as ESD, as evidenced by the darker diagonal values in Fig. A3. This observation reinforces the existence of a trade-off between unlearning effectiveness and retainability in the visual generative MU task, a phenomenon paralleled in other tasks such as classification.
D.3 Unlearning Heatmaps of More Unlearning Methods
[…]
@@ D.3 Unlearning Heatmaps of More Unlearning Methods
Figure A4: Heatmap visualization of FMN. The plot setting is identical to Figure 6.
Figure A5: Heatmap visualization of SEOT. The plot setting is identical to Figure 6.
Figure A6: Heatmap visualization of SPM. The plot setting is identical to Figure 6.
Figure A7: Heatmap visualization of Ediff. The plot setting is identical to Figure 6.
Figure A8: Heatmap visualization of SHS. The plot setting is identical to Figure 6.
E Visualizations
In this section, we aim to provide plenty of visualizations, illustrations, and qualitative results of the dataset and the quantitative results shown in Sec. 4 and Appx. D. These visualizations are intended to deepen the understanding of the effects of different MU methods and clearly illustrate the challenges identified in previous sections. We hope these visual aids will enable readers to more effectively grasp the nuances of MU methods and their implications for DMs (Diffusion Models).
Table A5: Performance comparison of different DM unlearning methods in the sequential unlearning setting. Each column represents a new unlearning request, denoted by $\mathcal{T}_{i}$, where $\mathcal{T}_{1}$ is the oldest. Each row represents the UA for a specific unlearning objective or the retaining accuracy (RA), given by the average of IRA and CRA. Results indicating unlearning rebound effect are highlighted in orange, and those signifying catastrophic retaining failure are marked in red.
Method: ESD Method: FMN
Metrics $\mathcal{T}_{1}$ $\mathcal{T}_{1}\sim\mathcal{T}_{2}$ $\mathcal{T}_{1}\sim\mathcal{T}_{3}$ $\mathcal{T}_{1}\sim\mathcal{T}_{4}$ $\mathcal{T}_{1}\sim\mathcal{T}_{5}$ $\mathcal{T}_{1}\sim\mathcal{T}_{6}$ Metrics $\mathcal{T}_{1}$ $\mathcal{T}_{1}\sim\mathcal{T}_{2}$ $\mathcal{T}_{1}\sim\mathcal{T}_{3}$ $\mathcal{T}_{1}\sim\mathcal{T}_{4}$ $\mathcal{T}_{1}\sim\mathcal{T}_{5}$ $\mathcal{T}_{1}\sim\mathcal{T}_{6}$
————————– Unlearning Request ————————– ————————– Unlearning Request ————————–
UA $\mathcal{T}_{1}$ $100\%$ $99\%$ $95\%$ $87\%$ $81\%$ $75\%$ UA $\mathcal{T}_{1}$ $88\%$ $99\%$ $99\%$ $98\%$ $99\%$ $99\%$
$\mathcal{T}_{2}$ $100\%$ $100\%$ $96\%$ $87\%$ $79\%$ $\mathcal{T}_{2}$ $95\%$ $99\%$ $99\%$ $98\%$ $99\%$
$\mathcal{T}_{3}$ $100\%$ $98\%$ $99\%$ $98\%$ $\mathcal{T}_{3}$ $97\%$ $98\%$ $99\%$ $99\%$
$\mathcal{T}_{4}$ $100\%$ $100\%$ $99\%$ $\mathcal{T}_{4}$ $99\%$ $99\%$ $99\%$
$\mathcal{T}_{5}$ $100\%$ $99\%$ $\mathcal{T}_{5}$ $99\%$ $99\%$
$\mathcal{T}_{6}$ $100\%$ $\mathcal{T}_{6}$ $100\%$
RA $77.46\%$ $52.94\%$ $35.99\%$ $24.86\%$ $18.69\%$ $12.95\%$ RA $82.39\%$ $14.56\%$ $13.34\%$ $10.42\%$ $9.83\%$ $8.76\%$
Method: UCE Method: CA
Metrics $\mathcal{T}_{1}$ $\mathcal{T}_{1}\sim\mathcal{T}_{2}$ $\mathcal{T}_{1}\sim\mathcal{T}_{3}$ $\mathcal{T}_{1}\sim\mathcal{T}_{4}$ $\mathcal{T}_{1}\sim\mathcal{T}_{5}$ $\mathcal{T}_{1}\sim\mathcal{T}_{6}$ Metrics $\mathcal{T}_{1}$ $\mathcal{T}_{1}\sim\mathcal{T}_{2}$ $\mathcal{T}_{1}\sim\mathcal{T}_{3}$ $\mathcal{T}_{1}\sim\mathcal{T}_{4}$ $\mathcal{T}_{1}\sim\mathcal{T}_{5}$ $\mathcal{T}_{1}\sim\mathcal{T}_{6}$
————————– Unlearning Request ————————– ————————– Unlearning Request ————————–
UA $\mathcal{T}_{1}$ $93\%$ $95\%$ $98\%$ $96\%$ $97\%$ $98\%$ UA $\mathcal{T}_{1}$ $58\%$ $55\%$ $59\%$ $45\%$ $44\%$ $40\%$
$\mathcal{T}_{2}$ $97\%$ $98\%$ $98\%$ $98\%$ $95\%$ $\mathcal{T}_{2}$ $76\%$ $58\%$ $51\%$ $47\%$ $44\%$
$\mathcal{T}_{3}$ $95\%$ $97\%$ $98\%$ $99\%$ $\mathcal{T}_{3}$ $45\%$ $41\%$ $40\%$ $37\%$
$\mathcal{T}_{4}$ $98\%$ $98\%$ $98\%$ $\mathcal{T}_{4}$ $71\%$ $70\%$ $60\%$
$\mathcal{T}_{5}$ $97\%$ $99\%$ $\mathcal{T}_{5}$ $69\%$ $51\%$
$\mathcal{T}_{6}$ $99\%$ $\mathcal{T}_{6}$ $57\%$
RA $81.42\%$ $29.38\%$ $18.72\%$ $15.34\%$ $13.32\%$ $11.31\%$ RA $97.24\%$ $93.39\%$ $84.46\%$ $79.32\%$ $71.40\%$ $60.53\%$
Method: SalUn Method: SPM
Metrics $\mathcal{T}_{1}$ $\mathcal{T}_{1}\sim\mathcal{T}_{2}$ $\mathcal{T}_{1}\sim\mathcal{T}_{3}$ $\mathcal{T}_{1}\sim\mathcal{T}_{4}$ $\mathcal{T}_{1}\sim\mathcal{T}_{5}$ $\mathcal{T}_{1}\sim\mathcal{T}_{6}$ Metrics $\mathcal{T}_{1}$ $\mathcal{T}_{1}\sim\mathcal{T}_{2}$ $\mathcal{T}_{1}\sim\mathcal{T}_{3}$ $\mathcal{T}_{1}\sim\mathcal{T}_{4}$ $\mathcal{T}_{1}\sim\mathcal{T}_{5}$ $\mathcal{T}_{1}\sim\mathcal{T}_{6}$
————————– Unlearning Request ————————– ————————– Unlearning Request ————————–
UA $\mathcal{T}_{1}$ $84\%$ $79\%$ $78\%$ $65\%$ $67\%$ $64\%$ UA $\mathcal{T}_{1}$ $55\%$ $59\%$ $50\%$ $49\%$ $47\%$ $48\%$
$\mathcal{T}_{2}$ $81.42\%$ $75\%$ $72\%$ $69\%$ $61\%$ $\mathcal{T}_{2}$ $62\%$ $59\%$ $58\%$ $60\%$ $63\%$
$\mathcal{T}_{3}$ $90\%$ $85\%$ $84\%$ $87\%$ $\mathcal{T}_{3}$ $42\%$ $39\%$ $40\%$ $41\%$
$\mathcal{T}_{4}$ $84\%$ $86\%$ $81\%$ $\mathcal{T}_{4}$ $57\%$ $59\%$ $60\%$
$\mathcal{T}_{5}$ $79\%$ $81\%$ $\mathcal{T}_{5}$ $51\%$ $51\%$
$\mathcal{T}_{6}$ $89\%$ $\mathcal{T}_{6}$ $43\%$
RA $85.43\%$ $80.32\%$ $71.42\%$ $65.41\%$ $63.24\%$ $60.19\%$ RA $72.39\%$ $70.42\%$ $67.89\%$ $60.45\%$ $55.32\%$ $51.12\%$
Method: EDiff Method: SHS
Metrics $\mathcal{T}_{1}$ $\mathcal{T}_{1}\sim\mathcal{T}_{2}$ $\mathcal{T}_{1}\sim\mathcal{T}_{3}$ $\mathcal{T}_{1}\sim\mathcal{T}_{4}$ $\mathcal{T}_{1}\sim\mathcal{T}_{5}$ $\mathcal{T}_{1}\sim\mathcal{T}_{6}$ Metrics $\mathcal{T}_{1}$ $\mathcal{T}_{1}\sim\mathcal{T}_{2}$ $\mathcal{T}_{1}\sim\mathcal{T}_{3}$ $\mathcal{T}_{1}\sim\mathcal{T}_{4}$ $\mathcal{T}_{1}\sim\mathcal{T}_{5}$ $\mathcal{T}_{1}\sim\mathcal{T}_{6}$
————————– Unlearning Request ————————– ————————– Unlearning Request ————————–
UA $\mathcal{T}_{1}$ $97\%$ $93\%$ $91\%$ $93\%$ $85\%$ $90\%$ UA $\mathcal{T}_{1}$ $81\%$ $73\%$ $74\%$ $93\%$ $94\%$ $97\%$
$\mathcal{T}_{2}$ $92\%$ $89\%$ $93\%$ $91\%$ $87\%$ $\mathcal{T}_{2}$ $69\%$ $61\%$ $89\%$ $94\%$ $97\%$
$\mathcal{T}_{3}$ $96\%$ $93\%$ $90\%$ $84\%$ $\mathcal{T}_{3}$ $75\%$ $92\%$ $96\%$ $90\%$
$\mathcal{T}_{4}$ $91\%$ $92\%$ $90.22\%$ $\mathcal{T}_{4}$ $91\%$ $95\%$ $97\%$
$\mathcal{T}_{5}$ $99\%$ $97\%$ $\mathcal{T}_{5}$ $92\%$ $96\%$
$\mathcal{T}_{6}$ $94\%$ $\mathcal{T}_{6}$ $94\%$
RA $92.34\%$ $89.37\%$ $14.35\%$ $12.31\%$ $12.82\%$ $7.42\%$ RA $88.41\%$ $84.32\%$ $73.98\%$ $69.19\%$ $10.76\%$ $10.11\%$
E.1 Illustrations of the Styles and Objects in UnlearnCanvas
We first provide an illustration of the styles and object classes included in UnlearnCanvas. Specifically, we show the styles in Fig. A9 and object classes in Fig. A10 with the style and object names disclosed in the captions.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(11) (12) (13) (14) (15) (16) (17) (18) (19) (20)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(21) (22) (23) (24) (25) (26) (27) (28) (29) (30)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(31) (32) (33) (34) (35) (36) (37) (38) (39) (40)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(41) (42) (43) (44) (45) (46) (47) (48) (49) (50)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(51) (52) (53) (54) (55) (56) (57) (58) (59) (60)
Figure A9: An illustration of the images in each style in UnlearnCanvas used in, which are stylized from the same seed image from the ‘Dogs’ object class. The seed image is presented in Fig. A10 (6). Images are cropped and down-scaled for illustration purpose. The name of the styles are: (1) Abstractionism; (2) Artist Sketch; (3) Blossom Season; (4) Blue Blooming; (5) Bricks; (6) Byzantine; (7) Cartoon; (8) Cold Warm; (9) Color Fantasy; (10) Comic Etch; (11) Crayon; (12) Crypto Punks; (13) Cubism; (14) Dadaism; (15) Dapple; (16) Defoliation; (17) Dreamweave; (18) Early Autumn; (19) Expressionism; (20) Fauvism; (21) Foliage Patchwork; (22) French; (23) Glowing Sunset; (24) Gorgeous Love; (25) Greenfield; (26) Impasto; (27) Impressionism; (28) Ink Art; (29) Joy; (30) Liquid Dreams; (31) Palette Knife; (32) Magic Cube; (33) Meta Physics; (34) Meteor Shower; (35) Monet; (36) Mosaic; (37) Neon Lines; (38) On Fire; (39) Pastel; (40) Pencil Drawing; (41) Picasso; (42) Pointillism; (43) Pop Art; (44) Rainwash; (45) Realistic Watercolor; (46) Red Blue Ink; (47) Rust; (48) Sketch; (49) Sponge Dabbed; (50) Structuralism; (51) Superstring; (52) Surrealism; (53) Techno; (54) Ukiyoe; (55) Van Gogh; (56) Vibrant Flow; (57) Warm Love; (58) Warm Smear; (59) Watercolor; (60) Winter.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(11) (12) (13) (14) (15) (16) (17) (18) (19) (20)
Figure A10: An illustration of the seed images in each object class in UnlearnCanvas. Images are cropped and down-scaled for illustration purpose. The name of the object classes are: (1) Architecture; (2) Bear; (3) Bird; (4) Butterfly; (5) Cat; (6) Dog; (7) Fish; (8) Flame; (9) Flowers; (10) Frog; (11) Horse; (12) Human; (13) Jellyfish; (14) Rabbits; (15) Sandwich; (16) Sea; (17) Statue; (18) Tower; (19) Tree; (20) Waterfalls.
E.2 Visualization of Style Unlearning
In Fig. A11, we provide abundant generation examples of all the 9 methods benchmarked in this work in a case study of unlearning the ‘Cartoon’ style. Both the successful and failure cases are demonstrated in the context of unlearning effectiveness, in-domain retainability, and cross-domain retainability.
Figure A11: visualization of the unlearning performance of different methods on the task of style unlearning. Three text prompt templates are used to evaluate the unlearning effectiveness, in-domain retainability, and cross-domain retainability of each method. Images with green frame denote desirable results, while the ones with red frame denote unlearning or retaining failures.
E.3 Visualization of the Unlearning Performance in the Presence of Adversarial Prompts
In Fig. A12, we provide visualizations for the effect of adversarial prompts. As revealed in Fig. 8, all the DM unlearning methods experience a significant drop in unlearning effectiveness when attacked by the adversarial prompt, enabled by UnlearnDiffAtk [59]. We provide the image generation in four unlearning cases (two for style unlearning and two for object unlearning), and show the images of the unlearning target successfully generated in the presence of adversarial prompts.
Figure A12: Visualization of the images generated by the unlearned DMs using different unlearning methods in the absence or presence of adversarial prompts.
F Broader Use Cases of UnlearnCanvas
@@
F.1 Benchmarking Style Transfer using UnlearnCanvas
@@
Style transfer is a long-standing topic and focuses on transferring the artistic style from one style image (also known as the reference image) $\mathbf{x}_{s}$ to a target content image $\mathbf{x}_{c}$. The most recent methods typically employ a neural network, denoted by ${\bm{\theta}}_{s}$, to extract style features and perform stylization in a single inference step, expressed as $\hat{\mathbf{x}}_{\mathrm{o}}=f_{{\bm{\theta}}_{s}}(\mathbf{x}_{s},\mathbf{x}_{c})$. Current state-of-the-art (SOTA) style transfer techniques exhibit remarkable generalization capabilities, successfully transferring styles not encountered during training and not requiring any further back-propagations. Figure 1 illustrates the pipeline of this task. Most existing literature [74, 75, 76, 77, 78, 79] utilize WikiArt [66] to provide different styles for training the stylization network ${\bm{\theta}}_{s}$. During the evaluation, the validation set of WikiArt will serve as the style (reference) images, together with the content images from the COCO dataset [80], to form a test bed for style transfer and style learning methods. However, such a evaluation scheme has some inherent limitations, which will be detailed below.
@@
Although the task of style transfer has been widely studied, its evaluations are still based on very limited quantitative or even sorely qualitative assessments, potentially leading to incomplete and inaccurate assessments [81]. Upon examining the evaluation pipelines of over $10$ state-of-the-art (SOTA) style transfer methods, three significant challenges are identified within the current widely accepted evaluation frameworks. $\bullet$ Challenge I (C1): The lack of the ground truth images for style similarity evaluation. Unlike other vision tasks like classification [82], detection [83], and segmentation [84], one of the key shortcomings of the current style transfer evaluation lies in the lack of the ground truth images $\mathbf{x}_{g}$ for the given reference style image $\mathbf{x}_{s}$ and the content image $\mathbf{x}_{c}$. Consequently, existing evaluation metrics, such as the style loss $\ell_{\mathrm{style}}$ [85, 86], has to be calculated with the reference style image $\mathbf{x}_{s}$ as the ground truth $\mathbf{x}_{g}$, namely $\ell_{\mathrm{style}}(\mathbf{x}_{s},\hat{\mathbf{x}}_{o})$, rather than directly using the ground truth $\ell_{\mathrm{style}}(\mathbf{x}_{o},{\mathbf{x}}_{g})$. Obviously, such a indirect evaluation may lead to inaccurate results due to the different contents held in $\mathbf{x}_{s}$ and $\mathbf{x}_{o}$. Existing work has demonstrated that such evaluation metrics can lead to very different conclusion from that of the user study [79]. Therefore, the creation of a supervised dataset with ground truth stylized images for every content image under each style is a timely remedy. $\bullet$ Challenge II (C2): The lack of algorithm stability evaluation against varied reference images $\mathbf{x}_{s}$. The evaluation of algorithm stability in style transfer has often been neglected. This assessment requires consistent performance of a method on a content image $\mathbf{x}_{c}$ across different style reference images ${\mathbf{x}_{s}}$ representing the same target artistic style. Ideally, an algorithm should maintain uniform quality across various references within a style, avoiding significant performance variations. Current challenges in such evaluations stem from the lack of reference sets that exhibit high stylistic consistency within each style. Figure A1 showcases examples from the widely-used WikiArt dataset. Despite sharing the same artistic label, images within a row show considerable divergence in visual appearance and style. Consequently, using these images as style references can lead to stable algorithms producing stylistically varied outputs, leading to misleading assessment results. Therefore, creating a dataset with high stylistic uniformity within each style category and clear differentiation between styles is crucial for accurate measurement of algorithm stability.
@@
To revolutionize the evaluation framework with UnlearnCanvas, it is crucial to first understand the components that form a comprehensive evaluation pipeline, ensuring a thorough and impartial assessment of performance. In the realm of style transfer, three pivotal metrics stand paramount, as depicted in Figure A14:
@@
① Style Loss: Utilizes a feature map-based style loss [87] to quantify the stylistic dissimilarity between image pairs, effectively representing the inverse of style authenticity. ② Content Loss: Employs a VGG-based, feature-map content loss [87, 88] to measure the visual dissimilarity between the reference and generated images, essentially mirroring content integrity. ③ Averaged Standard Deviation (STD): Computes the average STD of style and content loss w.r.t. the same reference image, reflecting algorithm stability.
@@
Following this evaluation pipeline, we scrutinized $9$ prominent style transfer methods, including SaNet [78], Mcc [89], Mast [90], ArtFlow with its two variants (AF-AdaIn and AF-WCT) [79], IE-Contrast [91], Cast [88], StyTr2 [87], and BLIP [92].
@@
F.2 Other Possible Applications of UnlearnCanvas